water has a mass per mole of 18.0 g/mol, and each water molecule (H2O) has 10 electrons. a) how many electrons are there in one liter (1.00 x 10^-2 m^3) of water? b) what is the net charge of all these electrons?
Answer #1

Here ,

a) for one liter of water ,

mass of water = 1 Kg = 1000 gm

number of moles = 1000/18

total number of electrons = number of moles * 10 * 6.022 *10^23

total number of electrons = 1000/18 * 10 * 6.022 *10^23

total number of electrons = 3.346 *10^26

the total numbner of electrons is 3.346 *10^26

part B)

total charge of all electrons = 3.346 *10^26 * 1.602 *10^-19

total charge of all electrons = 53602920 C